Work from a Constant Force

Learning Goal: To understand how to compute thework done by a constant force acting on a particle that moves in astraight line.

In this problem, you will calculate the work done by a constantforce. A force is considered constant if vec F(vec r) is independent of r_vec. This is the most frequently encountered situation inelementary Newtonian mechanics.MWE_wi_0_v1_a.jpgpart AConsider a particle moving in a straight linefrom initial point B to final point A, acted upon by a constantforce F_vec.View FigureThe force (think of it as a field,having a magnitude and direction atevery position r_vec) is indicated by a series ofidentical vectors pointingto the left, parallel to the horizontal axis. The vectors are allidentical only because the force is constant along the path.Themagnitude of the force is F,and the displacement vector frompoint B to point A is L_vec (of magnitude L,making and angle theta (radians) with the positive x axis). FindW_BA, the work that the force F_vec performs on the particle as it moves from point B topoint A.Express the work in terms ofL,F,and theta. Remember to use radians, not degrees, for anyangles that appear in youranswer.MWE_wi_0_v1_a.jpgMWE_wi_0_v1_b.jpgpart BNow consider the same force F_vec acting on aparticle that travels from point A to pointB.View FigureThe displacement vector L_vec now points in the opposite direction as it did in Part A.Find the work W_AB done by F_vec in this case.Express your answer in terms ofL,F,and theta.
Answer #1
The work done is given byW = F . L=FL Cos θ 'where θ ' = 1.745*10-2θ[ θ in degrees , θ ' -radians ]W = 1.745*10-2 FL Cosθb) W = - 1.745*10-2 FL Cosθwork is said to be negative as the direction isopposite.
answered by: sinthia